# Difference Decomposition Networks for Infrared Small Target Detection

### v1

@@ III-B Basis Decomposition Module
Figure 3: Overview of the BDM.
@@
where $B$ is the batch size, $F$ is the feature dimension, and $F_{inp}$ is the input feature dimension. $\mathrm{L2Norm}$ refers to L2 normalization to ensure each basis feature has the same influence on $\vec{y}$.
@@
where $c$ refers to the number of input channels, while $\cdot$$\cdot$] is the concatenation operator applied along the certain dimension. $G$ represents the number of feature groups, as we can regard several features across channels as one group.
@@
((a))
((b))
Figure 4: Detailed structure of the SD2M and SD3M used in Figures 6 and 8 respectively, are constructed based on the BDM.
Figure 5: Our backbone adopts a three-stage U-Module structure, where each U-Module is an encoder–decoder architecture. The first U-Module extracts coarse information, the second U-Module performs as the main feature-extracted stage to obtain the refine information, and the third U-Module conducts fine-grained enhancement.
@@ III-C Spatial Difference Decomposition Network
As illustrated in Figures 4(a) and 4(b), we adjust the BDM for ISTD and introduce two modules, named SD2M and SD3M, to enhance the targets and suppress the clutter. In addition, we add those modules to our backbone, as shown in Figure 5, and propose the SD2Net, as illustrated in 6.
@@ III-C1 Backbone
$O=\mathcal{U}_{3}\big(\mathcal{U}_{2}(\mathcal{U}_{1}(I))\big).$
Figure 7: Architecture of the DFEM, which employs 8 convolutional kernels with $n$ dilation rates to extract multi-scale basis features.
Figure 8: The architecture of Inception Pooling.
@@ III-C2 Spatial Difference Decomposition Module
where H, W, and C denote the height, width, and channel dimension. Difference Feature Extraction Module (DFEM) uses the difference kernels to extract the high-frequency as basis features ${P^{d_{j}}_{i}}\in\mathbb{R}^{B\times C\times 1\times(HW)}$ (where i and j index the kernel type and the dilation rate). These features are then used to form the tensor $P_{m}\in\mathbb{R}^{B\times C\times(8n)\times(HW)}$:
@@
Figure 9: The structure of STD2Net.
@@ III-D Spatiotemporal Difference Decomposition Network
Since background clutter exhibits random motion patterns, while targets move consistently. Therefore, we introduce TD2M to motion information and propose STD2Net. The basis features extracted by TD2M consist of both the differences in SD2Net’s results between the current frame and the reference frames, and the current frame’s SD2Net output. The former captures motion information, while the latter provides current state information.
@@ IV Experiments
We present the experiment details, validate modules via ablation studies, and demonstrate the performance of BDM’s adaptations through comparisons between SISTD and MISTD.
@@ IV-A1 Experiment Dataset
For the task of the SISTD, we employ two widely recognized public benchmarks: NUDT-SIRST [RenLiHanShu2021DNANet] and IRSTD-1K [Zhang2022ISNet]. To evaluate performance on the more complex task of MISTD, we benchmark our network against existing SOTA methods using the comprehensive NUDT-MIRSDT dataset [Li2025dtum]. To ensure a fair and consistent evaluation across all experiments, we divide each dataset into two parts: 80% of the images are randomly selected to constitute the training set, while the remaining 20% are reserved as the test set for the final performance assessment. All comparative methods were evaluated on the same dataset to ensure a fair comparison.
TABLE I: Impact of orthogonality and non-orthogonality.
@@
TABLE II: Impact of SD2Net’s channel configs on NUDT-SIRST dataset. We bold the best results and underline the second best.
@@
TABLE III: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.
@@ IV-A2 Implementation Details
The loss function for all models is the IoU loss [yu2016unitbox]. All experiments are conducted on Ubuntu 20.04 LTS. We train SISTD models on an NVIDIA GeForce RTX 4080 GPU for 400 epochs, using the Adam optimizer with a learning rate of $5\times 10^{-4}$. The batch size is set to 4. The learning rate is decayed to $5\times 10^{-5}$ at epoch 200 and further reduced to $5\times 10^{-6}$. Our MISTD architecture runs on two 24GB GeForce RTX 3090 GPUs. We train for 20 epochs with the Adam optimizer and an initial learning rate of 0.001. Furthermore, we set the temporal window size $N$ of STD2Net to 5. We run each algorithm three times and use the average outcome as the final result.
@@ IV-A3 Evaluation Metrics
We adopt four evaluation metrics: false alarm rate ($\mathrm{P_{a}}$), probability of detection ($\mathrm{P_{d}}$), mean intersection-over-union (mIoU), and area under curve (AUC) for accuracy-based evaluation. In addition, we also use the frames per second (FPS), giga floating-point operations per second (GFLOPS), and parameter size as our efficiency-based evaluation.
IV-B Ablation Study in SISTD
IV-B1 Impact of Orthogonality
We conducted an ablation study to show the effects of using orthogonal versus non-orthogonal basis features, as summarized in Table I. The orthogonal basis features were obtained by applying singular value decomposition (SVD) to the original features. Experiments reveal that orthogonal features do not lead to any improvement in accuracy within our modules. On the contrary, they introduce a significant computational overhead. Consequently, the orthogonalization is unnecessary for our features.
IV-B2 Impact of the Number of Channels
To optimize channel dimension, we conduct a study on channel configurations, as detailed in Table II. We scale channel configurations from [4, 8, 16, 32] to [32, 64, 128, 256], yielding consistent improvements in mIoU. However, this comes at an increasing computational cost. The [8, 16, 32, 64] configuration achieves the optimal balance of performance across all metrics, and we adopt it as our channel configuration.
IV-B3 Module-wise Metrics Performance
We conduct ablation studies to validate the effectiveness of each module in our framework. The experimental results are summarized in Table III. All experiments are conducted on two SISTD datasets: NUDT-SIRST and IRSTD-1K. Using only the SD2M module achieves competitive results on NUDT-SIRST, confirming its effectiveness for target feature extraction. Additionally, the InceptionPooling module and SD3M module also enhance the network’s performance. Integrating all three modules achieves the optimal balance of performance.
TABLE IV: Quantitative comparison with different decomposition approaches on NUDT-SIRST and IRSTD-1K.
@@ IV-B3 Module-wise Metrics Performance
IV-C Comparison with State-of-the-art (SOTA) Methods in SISTD
TABLE V: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.

### v2

@@ III-B1 Overall Procedure
((a))
((b))
Figure 5: Detailed structure of the SD2M and SD3M used in Figures 7 and 9 respectively, are constructed based on the BDM. SD2M is utilized for the feature extraction without
Figure 6: Our backbone adopts a three-stage U-Module structure, where each U-Module is an encoder–decoder architecture. The first U-Module extracts coarse information, the second U-Module performs as the main feature-extraction stage to obtain the refined information, and the third U-Module conducts fine-grained enhancement.
@@
where $B$ is the batch size, $F$ is the feature dimension, and $F_{inp}$ is the input feature dimension. $G$ represents the number of feature groups, as we can regard several features across channels as one group. $\mathrm{L2Norm}$ refers to L2 normalization to ensure each basis feature has the same influence on $\vec{y}$.
@@
where $c$ refers to the number of input channels, while $\cdot$$\cdot$] is the concatenation operator applied along the certain dimension.
Figure 8: Architecture of the DFEM, which employs eight difference convolution kernels with $n$ dilation rates to extract multi-scale basis features.
@@
III-B2 Difference Decomposition Module
The primary difference between the proposed DDMs and the original module lies in the use of the BFEM. This module dynamically derives difference features across spatial and temporal dimensions using a variety of techniques, including central difference convolutions and the difference between two frames.
@@ III-C Spatial Difference Decomposition Network
As illustrated in Figures 5(a) and 5(b), we adjust the BDM for IRSTD and introduce two DDMs, named Spatial Difference Decomposition Module (SD2M) and Spatial Difference Decomposition Downsampling Module (SD3M), to enhance the targets and suppress clutter. In addition, we add those modules to our backbone, as shown in Figure 6, and propose the SD2Net, as illustrated in 7.
SD2M serves as a difference information extractor that maintains resolution. SD3M acts as a detail-preserving downsampling module, embedded into a multi-branch module.
@@ III-C1 Backbone
$O=\mathcal{U}_{3}\big(\mathcal{U}_{2}(\mathcal{U}_{1}(I))\big).$ (10)
@@ III-C2 Spatial Difference Decomposition Module
where H, W, and C denote the height, width, and channel dimension. The Difference Feature Extraction Module (DFEM) utilizes difference kernels to extract high-frequency basis features ${P^{d_{j}}_{i}}\in\mathbb{R}^{B\times C\times 1\times(HW)}$ (where i and j index the kernel type and dilation rate, respectively). These features are then used to form the tensor $P_{m}\in\mathbb{R}^{B\times C\times(8n)\times(HW)}$:
@@
Figure 9: The architecture of Inception Pooling.
Figure 10: The structure of STD2Net comprises two main components: the TFEM is employed to extract difference basis features in the temporal domain. At the same time, the FEM is utilized to fuse the reference frames with the key frame to obtain the features with temporal information.
@@
While SD3M could keep detailed information, the downsampling process itself requires broader contextual information beyond these details. Therefore, we utilize multiple branches to maintain various information.
@@ III-D Spatiotemporal Difference Decomposition Network
Since background clutter exhibits random motion patterns, while targets move consistently. Therefore, we introduce TD2M to motion information and propose STD2Net, as shown in Figure 10. The basis features extracted by Temporal Difference Decomposition Module (TD2M) consist of both the differences in SD2Net’s results between the current frame and the reference frames, and the current frame’s SD2Net output. The former captures motion information, while the latter provides current state information.
@@ IV Experiments
We present the experiment details, validate modules via ablation studies, and demonstrate the performance of BDM’s adaptations through comparisons between SIRSTD and MIRSTD.
@@ IV-A1 Experiment Dataset
For the task of the SIRSTD, we employ two widely recognized public benchmarks: NUDT-SIRST [RenLiHanShu2021DNANet] and IRSTD-1K [Zhang2022ISNet]. To evaluate performance on the more complex task of MIRSTD, we benchmark our network against existing SOTA methods using the comprehensive NUDT-MIRSDT dataset [Li2025dtum]. To ensure a fair and consistent evaluation across all experiments, we divide each dataset into two parts: 80% of the images are randomly selected to constitute the training set, while the remaining 20% are reserved as the test set for the final performance assessment. All comparative methods were evaluated on the same dataset to ensure a fair comparison.
TABLE I: Impact of orthogonality and normalization.
@@
TABLE II: Impact of SD2Net’s dilation configs on NUDT-SIRST dataset. We bold the best results and underline the second best.
@@
TABLE III: Impact of SD2Net’s channel configs on NUDT-SIRST dataset. We bold the best results and underline the second best.
[…]
TABLE IV: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.
@@ IV-A2 Implementation Details
The loss function for all models is the IoU loss [yu2016unitbox]. All experiments are conducted on Ubuntu 20.04 LTS. We train SIRSTD models on an NVIDIA GeForce RTX 4080 GPU for 400 epochs, using the Adam optimizer with a learning rate of $5\times 10^{-4}$. The batch size is set to 4. The learning rate is decayed to $5\times 10^{-5}$ at epoch 200 and further reduced to $5\times 10^{-6}$. Our MIRSTD architecture runs on two 24GB GeForce RTX 3090 GPUs. We train for 20 epochs with the Adam optimizer and an initial learning rate of 0.001. Furthermore, we set the temporal window size $N$ of STD2Net to 5. We run each algorithm three times and use the average outcome as the final result. All models, including the baselines, are trained under our experimental configuration to ensure a fair comparison.
@@ IV-A3 Evaluation Metrics
We adopt four evaluation metrics: false alarm rate ($\mathrm{P_{a}}$), probability of detection ($\mathrm{P_{d}}$), mean intersection-over-union (mIoU), and area under curve (AUC) for accuracy-based evaluation. In addition, we also use frames per second (FPS), giga floating-point operations per second (GFLOPS), and parameter size as our efficiency-based evaluation metrics.
IV-B Ablation Study in SIRSTD
IV-B1 Impact of Orthogonality and Normalization
We conducted an ablation study to show the effects of using orthogonal versus non-orthogonal basis features, as summarized in Table I. The orthogonal basis features were obtained by applying singular value decomposition (SVD) to the original features. Experiments reveal that orthogonal features do not lead to any improvement in accuracy within our modules. On the contrary, they introduce a significant computational overhead. However, the normalization is the key part of our method. Consequently, we retain the normalization and relax the orthogonality constraint.
IV-B2 Impact of the Dilation Ratios
To determine the optimal setting, we performed an ablation study on dilation ratios, systematically evaluating the performance of various configurations, as shown in Table II. The experiment results indicated that the [1, 3] configuration significantly outperformed the others. Based on these findings, we adopted [1, 3] as our dilation configuration. We apply a dilation ratio of 3 to larger targets and a ratio of 1 to smaller targets.
IV-B3 Impact of the Number of Channels
To optimize channel dimension, we conduct a study on channel configurations, as detailed in Table III. We scale channel configurations from [4, 8, 16, 32] to [32, 64, 128, 256], yielding consistent improvements in mIoU. However, this comes at an increasing computational cost. The [8, 16, 32, 64] configuration achieves the optimal balance of performance across all metrics, and we adopt it as our channel configuration.
IV-B4 Module-wise Metrics Performance
The SD2M works for enhancing the high-frequency targets. Additionally, the SD3M and InceptionPool are utilized for downsampling, which removes less task-relevant information. We conduct ablation studies to validate the effectiveness of each module in our framework. The experimental results are summarized in Table IV. All experiments are conducted on two SIRSTD datasets: NUDT-SIRST and IRSTD-1K. Using only the SD2M module achieves competitive results on NUDT-SIRST, confirming its effectiveness for target feature extraction. Additionally, the InceptionPooling module and SD3M module also enhance the network’s performance. Integrating all three modules achieves the optimal balance of performance. Therefore, all modules are effective.
TABLE V: Quantitative comparison with different decomposition approaches on NUDT-SIRST and IRSTD-1K.
@@ IV-B4 Module-wise Metrics Performance
IV-C Comparison with State-of-the-art (SOTA) Methods in SIRSTD
TABLE VI: Ablation studies of SD2Net using NUDT-SIRST and IRSTD-1K. We bold the best results and underline the second best.